# Topological constraints on clean Lagrangian intersections from $\mathbb{Q}$-valued augmentations

### v2

@@ 3.2.1 Review of the definition of $HC_{0}(K)$
First, let us review how to compute $HC_{0}(K)$ for an oriented knot $K$. We refer to [15, Section 3]. Note that since $\mathcal{A}^{\mathrm{knot}}_{p}(K)=0$ for $p<0$, $HC_{0}(K)=\mathcal{A}^{\mathrm{knot}}_{0}(K)/\partial_{K}(\mathcal{A}^{\mathrm{knot}}_{1}(K))$.
For any $n\in\mathbb{Z}_{\geq 1}$, let $B_{n}$ denote the group of braids on $n$ strands, which is generated by $\{\sigma_{1},\dots,\sigma_{n-1}\}$ with relations $\sigma_{i}\sigma_{i+1}\sigma_{i}=\sigma_{i+1}\sigma_{i}\sigma_{i+1}$ for $1\leq i\leq n-2$ and $\sigma_{i}\sigma_{j}=\sigma_{j}\sigma_{i}$ for $|i-j|\geq 2$. Let $\mathcal{A}_{n}$ be the $R$-algebra freely generated by the set $\{a_{i,j}\mid i,j\in\{1,\dots,n\},\ i\neq j\}$. Then, a group homomorphism $B_{n}\to\mathrm{Aut}(\mathcal{A}_{n})\colon B\mapsto\phi_{B}$ is defined as in [15, Definition 3.3]. It is determined by $\phi_{\sigma_{k}}\colon\mathcal{A}_{n}\to\mathcal{A}_{n}$ for each generator $\sigma_{k}$, and $\phi_{\sigma_{k}}$ is given by
@@
The automorphism $\phi_{B}\in\mathrm{Aut}(\mathcal{A}_{n})$ for $B\in B_{n}$ is extended to an automorphism on $\mathcal{A}_{n+1}$ as follows: We consider the group homomorphism $B_{n+1}\to\mathrm{Aut}(\mathcal{A}_{n+1})$ defined as above and an injective map $B_{n}\to B_{n+1}$ defined by adding the extra $(n+1)$-th strand to any braid on $n$ strands (see [15, Remark 3.5]). Then, for any $B\in B_{n}$, $\phi_{B}\in\mathrm{Aut}(\mathcal{A}_{n+1})$ is defined via the composite map $B_{n}\hookrightarrow B_{n+1}\to\mathrm{Aut}(\mathcal{A}_{n+1})$. In addition, let us label the $(n+1)$-th strand by * and denote $a_{i,n+1}$ by $a_{i,*}$ and $a_{n+1,i}$ by $a_{*,i}$ for every $i\in\{1,\dots,n\}$.
@@
Let $K$ be the oriented knot in $\mathbb{R}^{3}$ given by the closure of the braid $B$. By [15, Theorem 4.1, Remark 4.2], $HC_{0}(K)$ is the quotient of $\mathcal{A}_{n}$ by the two-sided ideal generated by all entries of the two matrices
@@
where $(\phi_{B}(\mathbb{A}))_{i,j}\coloneqq\phi_{B}(\mathbb{A}_{i,j})$ for $i,j\in\{1,\dots,n\}$. The isomorphism class of $HC_{0}(K)$ as a unital $R$-algebra depends only on the isotopy class of $K$ as an oriented knot. See [15, Theorem 3.15] and [14, Theorem 1.4 (2)].
@@ 3.2.2 On augmentations of $\mathcal{A}^{\mathrm{knot}}_{*}(K\#T_{(2,2m+1)})$ for $m\geq 0$
for every $m\in\mathbb{Z}_{\geq 0}$. Then, the closure of $B^{\#}_{m}$ gives the connected sum $K\#T_{(2,2m+1)}$ of $K$ and the $(2,2m+1)$-torus knot. See Figure 1 in the case of $m=1$. (When $m=0$, $T_{(2,1)}$ is the unknot.)
[…]
Figure. 1: The braid $B^{\#}_{m}\in B_{n+2}$ for $m=1$. The box with ‘$B$’ in the center represents the braid $B$ on $n$ strands.
@@
Let $\mathbb{k}$ be a field. For any augmentation $\varepsilon$ of $\mathcal{A}^{\mathrm{knot}}_{*}(K\#T_{(2,2m+1)})$ to $\mathbb{k}$ such that $(\varepsilon(\mu),\varepsilon(U))=(y,Z)\in(\mathbb{k}^{*})^{2}$, the polynomial $\left.P_{m}\right|_{\mu=y,U=Z}\in\mathbb{k}[T]$ has a root in $\mathbb{k}$.
@@
which maps $T$ to $XY=a_{n,n+2}a_{n+2,n}$. Now the proposition follows from Lemma 3.4. ∎
@@
From a braid representation $(\sigma_{1})^{3}\in B_{2}$ of $T_{(2,3)}$, it is not difficult to compute $HC_{0}(T_{(2,3)})$ as outlined in [15, Exercise 3.10]. $HC_{0}(T_{(2,3)})$ is isomorphic to a commutative $R$-algebra $R[X]/(G_{1},G_{2})$, where
@@
See [15, Exercise 5.5]. We have identities
@@
$UX(XG_{1}+\mu G_{2})+\lambda\mu^{2}(\mu-U)G_{1}$
$=\$$(UX^{2})^{2}+\lambda\mu^{2}(3\mu-\mu^{2}-U)(UX^{2})+\lambda^{2}\mu^{5}(1-\mu)(\mu-U)$
@@
Consider the reduction of the coefficient ring from $R$ to $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$ by setting $U=1$. By the substitution $U=1$, we have $\left.P_{0}\right|_{U=1}=(1-\mu)^{2}+\mu T$, and if we define $h^{\prime}_{m},k^{\prime}_{m}\in\mathbb{Z}[\mu^{\pm}]$ by $h^{\prime}_{m}\coloneqq\left.h_{m}\right|_{T=-\mu^{-1}(1-\mu)^{2}}$ and $k^{\prime}_{m}\coloneqq\left.k_{m}\right|_{T=-\mu^{-1}(1-\mu)^{2}}$, then
@@
Here, the second equality follows from (20). By induction on $m=0,1,\dots$, we can see that the polynomial $\left.P_{m}\right|_{U=1}\in\mathbb{Z}[\mu^{\pm}][T]$ has a root $T=-\mu^{-1}(1-\mu)^{2}$. Hence, Proposition 3.10 is trivial for if an augmentation $\varepsilon$ satisfies $\varepsilon(U)=1$.
@@ 3.2.3 Application of $\mathbb{Q}$-valued augmentations
Combining Proposition 3.5 and Proposition 3.10 with $\mathbb{k}=\mathbb{Q}$, we obtain the following theorem.
@@
For a knot $K_{0}$, suppose that there exists $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection $K_{1}=\varphi(L_{K_{0}})\cap\mathbb{R}^{3}$ and $K_{1}$ is the unknot in $\mathbb{R}^{3}$. Assuming that $K_{0}=K\#T_{(2,2m+1)}$, we will deduce contradictions in two cases: the case $m\geq 1$ and the case $m\leq-2$.
(The case $m\geq 1$) By Proposition 3.5, for every $Z_{0}\in\{r\in\mathbb{Q}^{*}\mid r^{n}\neq y_{0}\text{ for every }n\in\mathbb{Z}\}$, there exists an augmentation $\varepsilon$ of $(\mathcal{A}^{\mathrm{knot}}_{*}(K_{0}),\partial_{K_{0}})$ to $\mathbb{Q}$ such that $\varepsilon(\mu)=y_{0}$ and $\varepsilon(U)=Z_{0}$. Then, it follows from Proposition 3.10 that $\left.P_{m}\right|_{\mu=y_{0},U=Z_{0}}\in\mathbb{Q}[T]$ has a root in $\mathbb{Q}$. From the definition (23) of $P_{m}$, it is equivalent to that the equation for $T$
$(Z_{0}-y_{0})\cdot\bar{h}_{m}(T)+y_{0}T\cdot\bar{k}_{m}(T)=0$ (24)
has a solution in $\mathbb{Q}$.
As a polynomial in $\mathbb{Q}(Z)[T]$, let us consider $(Z-y_{0})\cdot\bar{h}_{m}+y_{0}T\cdot\bar{k}_{m}$. Let $p_{1},\dots,p_{e}\in\mathbb{Q}(Z)[T]$ be irreducible elements such that
$(Z-y_{0})\cdot\bar{h}_{m}+y_{0}T\cdot\bar{k}_{m}=p_{1}\cdots p_{e}.$
@@
By Hilbert’s irreduciblity theorem (see e.g. [11, Chapter 9, Corollary 2.4], which shows that $\mathbb{Q}$ is a Hilbertian field), there are infinitely many $Z_{0}\in\mathbb{Q}$ such that all of the specializations
@@
are irreducible in $\mathbb{Q}[T]$. We note that the set $\bigcup_{i=1}^{e}\{Z_{0}\in\mathbb{Q}\mid\deg(\left.p_{i}\right|_{Z=Z_{0}})\leq 1\}$ is finite from the above observation of the degree of $p_{1},\dots,p_{m}$ in $T$. In addition, for the fixed rational number $y_{0}\neq 1$, the set $\{r\in\mathbb{Q}^{*}\mid r^{n}=y_{0}\text{ for some }n\in\mathbb{Z}\}$ is also finite. As a consequence, there are infinitely many $Z_{0}\in\mathbb{Q}^{*}\setminus\{r\in\mathbb{Q}^{*}\mid r^{n}=y_{0}\text{ for some }n\in\mathbb{Z}\}$ such that $\left.p_{i}\right|_{Z=Z_{0}}\in\mathbb{Q}[T]$ is irreducible and $\deg(\left.p_{i}\right|_{Z=Z_{0}})\geq 2$ for every $i\in\{1,\dots,e\}$, which means that the equation (24) has no solution in $\mathbb{Q}$. This contradicts Proposition 3.5 and Proposition 3.10.
(The case $m\leq-2$) Take a Hamiltonian $H\colon T^{*}\mathbb{R}^{3}\times[0,1]\to\mathbb{R}$ such that $\varphi=\varphi^{1}_{H}$. Consider an involution
@@
$\varphi^{1}_{\overline{H}}(L_{\overline{K_{0}}})$ intersects $\mathbb{R}^{3}=\iota(\mathbb{R}^{3})$ cleanly along $\overline{K_{1}}$. Since $\overline{K_{1}}$ is the unknot and $\overline{K_{0}}$ is the connected sum of the mirror of $K$ and $T_{(2,-2m-1)}$, this contradicts to the result from the above case. ∎
Remark 3.16.
To the author’s knowledge, it seems difficult to show Theorem 3.15, even when $K_{0}=T_{(2,3)}$, by only using the knot DGA with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$. Indeed, for the trefoil knot $T_{(2,3)}$ and the unknot $O$,
$\left.HC_{0}(T_{(2,3)})\right|_{U=1}$ $\cong\mathbb{Z}[\lambda^{\pm},\mu^{\pm}][X]/(X^{2}-\mu X+\lambda\mu^{3}(1-\mu),\ X^{2}+\lambda\mu^{2}X+\lambda\mu^{2}(\mu-1)),$
$\left.HC_{0}(O)\right|_{U=1}$ $\cong\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]/((1-\lambda)(1-\mu)),$
and we can find a unital $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$-algebra map
$\left.HC_{0}(T_{(2,3)})\right|_{U=1}\to\left.HC_{0}(O)\right|_{U=1}\colon X\mapsto\mu-\mu^{2}.$
Therefore, we cannot deduce a contradiction from the existence of a unital $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$-algebra map between the 0-th homology which is induced from the DGA map as in Theorem 3.2.
@@
Let $\mathbb{k}$ be a field. Then, for any augmentation $\varepsilon$ of $\mathcal{A}^{\mathrm{knot}}_{*}(K\#4_{1})$ to $\mathbb{k}$ such that $(\varepsilon(\mu),\varepsilon(U))=(y,Z)$, the polynomial $\left.P\right|_{\mu=y,U=Z}\in\mathbb{k}[T]$ has a root in $\mathbb{k}$.
@@
If we define $P\in\mathbb{Z}[\mu^{\pm},U][T]$ by (25), then
$(U-\mu)\cdot F+\mu Y\cdot G=P(XY),$
which implies that $P(XY)$ is contained in $\mathcal{I}_{B^{\#}}$. Therefore, we have a well-defined unital $R$-algebra map
@@
which maps $T$ to $XY=a_{n,n+3}a_{n+3,n}$. Now, the proposition follows from Lemma 3.4. ∎
@@
For a knot $K_{0}$, suppose that there exists $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection $K_{1}=\varphi(L_{K_{0}})\cap\mathbb{R}^{3}$ and $K_{1}$ is the unknot in $\mathbb{R}^{3}$. Assuming that $K_{0}=K\#4_{1}$, we deduce a contradiction as follows: By Proposition 3.5, there exists an augmentation $\varepsilon$ of $(\mathcal{A}^{\mathrm{knot}}_{*}(K_{0}),\partial_{K_{0}})$ to $\mathbb{Q}$ such that $\varepsilon(\mu)=-1$ and $\varepsilon(U)=-2$. Then, it follows from Proposition 3.17 that the polynomial $\bar{P}\coloneqq\left.P\right|_{\mu=-1,U=-2}\in\mathbb{Q}[T]$ has a root in $\mathbb{Q}$. On the other hand, from the definition of $P$,
$\bar{P}(T)=2-\frac{3}{2}T-2T^{3}$
Multiplying by $2$, we obtain a polynomial with integer coefficients
$4-3T-4T^{3}.$
By the rational root theorem, the only possible rational roots of this polynomial are of the form $\frac{p}{q}$, where $p\in\mathbb{Z}$ is a divisor of $4$ and $q\in\mathbb{Z}$ is a divisor of 4-4. The set of such rational numbers consists of the integers
$\pm 1,\pm 2,\pm 4,$
and those which are not integers
$\pm\frac{1}{2},\pm\frac{1}{4}.$
We note that $2\cdot\bar{P}(T)=4-T(4T^{2}+3)$. For every $n\in\{\pm 1,\pm 2,\pm 4\}$, the integer $4n^{2}+3$ is not a divisor of $4$, which implies that $2\cdot\bar{P}(n)\neq 0$. One can also check that
@@
Therefore, we can conclude that $\bar{P}$ does not have a root in $\mathbb{Q}$, which contradicts Proposition 3.5 and Proposition 3.17. ∎
@@
For any $\psi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$, we introduce a $\mathbb{Z}/2$-graded Floer cochain complex
@@
with coefficients in the Laurent polynomial ring $\mathbb{F}_{2}[\lambda^{\pm}]$. It is a slight enhancement of the Floer cochain complex over $\mathbb{F}_{2}$ constructed as in [19, Section 8].
We choose a Hamiltonian $H\colon T^{*}\mathbb{R}^{3}\times[0,1]\to\mathbb{R}$ with compact support such that the Lagrangian submanifold $P_{H}\coloneqq(\varphi^{1}_{H}\circ\psi)(\mathbb{R}^{3})$ intersects $L_{K_{0}}$ transversely. $P_{H}$ is oriented so that the diffeomorphism $\varphi^{1}_{H}\circ\psi\colon\mathbb{R}^{3}\to P_{H}$ preserves orientations. From the orientations of $P_{H}$ and $L_{K_{0}}$, a $\mathbb{Z}/2$-grading $|x|\in\mathbb{Z}/2$ is defined for every $x\in P_{H}\cap L_{K_{0}}$. See [18, Example 2.7, Section 2.d]. Note that the grading is shifted by $1$ if we use the opposite orientation of $P_{H}$. We also choose an almost complex structure $J$ such that $d\lambda_{\mathbb{R}^{3}}(\cdot,J\cdot)$ is a Riemannian metric on $T^{*}\mathbb{R}^{3}$ and $J(\partial_{p_{i}})=\partial_{q_{i}}$ for $i\in\{1,2,3\}$ outside a compact subset of $T^{*}\mathbb{R}^{3}$.
[…]

### v3

@@ 3.2.1 Review of the definition of $HC_{0}(K)$
First, let us review how to compute $HC_{0}(K)$ for an oriented knot $K$. We refer to [14, Section 3]. Note that since $\mathcal{A}^{\mathrm{knot}}_{p}(K)=0$ for $p<0$, $HC_{0}(K)=\mathcal{A}^{\mathrm{knot}}_{0}(K)/\partial_{K}(\mathcal{A}^{\mathrm{knot}}_{1}(K))$.
For any $n\in\mathbb{Z}_{\geq 1}$, let $B_{n}$ denote the group of braids on $n$ strands, which is generated by $\{\sigma_{1},\dots,\sigma_{n-1}\}$ with relations $\sigma_{i}\sigma_{i+1}\sigma_{i}=\sigma_{i+1}\sigma_{i}\sigma_{i+1}$ for $1\leq i\leq n-2$ and $\sigma_{i}\sigma_{j}=\sigma_{j}\sigma_{i}$ for $|i-j|\geq 2$. Let $\mathcal{A}_{n}$ be the $R$-algebra freely generated by the set $\{a_{i,j}\mid i,j\in\{1,\dots,n\},\ i\neq j\}$. Then, a group homomorphism $B_{n}\to\mathrm{Aut}(\mathcal{A}_{n})\colon B\mapsto\phi_{B}$ is defined as in [14, Definition 3.3]. It is determined by $\phi_{\sigma_{k}}\colon\mathcal{A}_{n}\to\mathcal{A}_{n}$ for each generator $\sigma_{k}$, and $\phi_{\sigma_{k}}$ is given by
@@
The automorphism $\phi_{B}\in\mathrm{Aut}(\mathcal{A}_{n})$ for $B\in B_{n}$ is extended to an automorphism on $\mathcal{A}_{n+1}$ as follows: We consider the group homomorphism $B_{n+1}\to\mathrm{Aut}(\mathcal{A}_{n+1})$ defined as above and an injective map $B_{n}\to B_{n+1}$ defined by adding the extra $(n+1)$-th strand to any braid on $n$ strands (see [14, Remark 3.5]). Then, for any $B\in B_{n}$, $\phi_{B}\in\mathrm{Aut}(\mathcal{A}_{n+1})$ is defined via the composite map $B_{n}\hookrightarrow B_{n+1}\to\mathrm{Aut}(\mathcal{A}_{n+1})$. In addition, let us label the $(n+1)$-th strand by * and denote $a_{i,n+1}$ by $a_{i,*}$ and $a_{n+1,i}$ by $a_{*,i}$ for every $i\in\{1,\dots,n\}$.
@@
Let $K$ be the oriented knot in $\mathbb{R}^{3}$ given by the closure of the braid $B$. By [14, Theorem 4.1, Remark 4.2], $HC_{0}(K)$ is the quotient of $\mathcal{A}_{n}$ by the two-sided ideal generated by all entries of the two matrices
@@
where $(\phi_{B}(\mathbb{A}))_{i,j}\coloneqq\phi_{B}(\mathbb{A}_{i,j})$ for $i,j\in\{1,\dots,n\}$. The isomorphism class of $HC_{0}(K)$ as a unital $R$-algebra depends only on the isotopy class of $K$ as an oriented knot. See [14, Theorem 3.15] and [13, Theorem 1.4 (2)].
@@ 3.2.2 On augmentations of $\mathcal{A}^{\mathrm{knot}}_{*}(K\#T_{(2,2m+1)})$ for $m\geq 0$
for every $m\in\mathbb{Z}_{\geq 0}$. Then, the closure of $B^{\#}_{m}$ gives the connected sum $K\#T_{(2,2m+1)}$ of $K$ and the $(2,2m+1)$-torus knot. (When $m=0$, $T_{(2,1)}$ is the unknot.)
@@
For any field $\mathbb{k}$,
$\pi_{y,Z}(V_{\mathbb{k}}(K\#T_{(2,2m+1)}))\subset\{(y,Z)\in(\mathbb{k}^{*})^{2}\mid\left.P_{m}\right|_{\mu=y,U=Z}\text{ has a root in }\mathbb{k}\}.$
@@
which maps $T$ to $XY=a_{n,n+2}a_{n+2,n}$.
Given a $\mathbb{k}$-valued augmentation $\varepsilon$ of $\mathcal{A}^{\mathrm{knot}}_{*}(K\#T_{(2,2m+1)})$, it induces a unital ring homomorphism from $HC^{\mathrm{ab}}_{0}(K\#T_{(2,2m+1)})$ to $\mathbb{k}$. By precomposing $J_{m}$, we obtain a unital ring homomorphism $R[T]/(P_{m})\to\mathbb{k}$ which sends $\mu$ to $\varepsilon(\mu)$ and $U$ to $\varepsilon(U)$. This means that $\left.P_{m}\right|_{\mu=\varepsilon(\mu),U=\varepsilon(U)}\in\mathbb{k}[T]$ has a root in $\mathbb{k}$. Now, the proposition follows from the definition of $V_{\mathbb{k}}(K\#T_{(2,2m+1)})$. ∎
@@
From a braid representation $(\sigma_{1})^{3}\in B_{2}$ of $T_{(2,3)}$, it is not difficult to compute $HC_{0}(T_{(2,3)})$ as outlined in [14, Exercise 3.10]. $HC_{0}(T_{(2,3)})$ is isomorphic to a commutative $R$-algebra $R[X]/(G_{1},G_{2})$, where
@@
See [14, Exercise 5.5]. We have identities
@@
$UX(XG_{1}+\mu G_{2})+\lambda\mu^{2}(\mu-U)G_{1}=\$$(UX^{2})^{2}+\lambda\mu^{2}(3\mu-\mu^{2}-U)(UX^{2})+\lambda^{2}\mu^{5}(1-\mu)(\mu-U)$
@@
Consider the reduction of the coefficient ring from $R$ to $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$ by setting $U=1$. In this case, we obtain $V_{\mathbb{k}}\cap\{Z=1\}\subset\{(x,y)\mid\left.P_{m}\right|_{\mu=y,U=1}\text{ has a root in }\mathbb{k}\}$. However, this is a trivial. By the substitution $U=1$, we have $\left.P_{0}\right|_{U=1}=(1-\mu)^{2}+\mu T$, and if we define $h^{\prime}_{m},k^{\prime}_{m}\in\mathbb{Z}[\mu^{\pm}]$ by $h^{\prime}_{m}\coloneqq\left.h_{m}\right|_{T=-\mu^{-1}(1-\mu)^{2}}$ and $k^{\prime}_{m}\coloneqq\left.k_{m}\right|_{T=-\mu^{-1}(1-\mu)^{2}}$, then
@@
Here, the second equality follows from (21). By induction on $m=0,1,\dots$, we can see that the polynomial $\left.P_{m}\right|_{U=1}\in\mathbb{Z}[\mu^{\pm}][T]$ has a root $T=-\mu^{-1}(1-\mu)^{2}$.
@@ 3.2.3 Application of $\mathbb{Q}$-valued augmentations
Now, we prove Theorem 1.1 for $K^{\prime}\#T_{(2,q)}$, where $q\neq\pm 1$ is an odd number.
@@
Let us first consider the case $m\geq 1$. From Proposition 3.6 and Proposition 3.9 with $\mathbb{k}=\mathbb{Q}$, it suffices to show that in $(\mathbb{Q}^{*})^{2}$
[…]
For the fixed rational number $y_{0}\neq 1$, note that
$\left.P_{m}\right|_{\mu=y_{0},U=Z}=(Z-y_{0})\cdot\bar{h}_{m}(T)+y_{0}T\cdot\bar{k}_{m}(T)$
for any $Z\in(\mathbb{Q}^{*})^{2}$. Let us regard $\left.P_{m}\right|_{\mu=y_{0},U=Z}$ as a polynomial in $\mathbb{Q}(Z)[T]$. Take $p_{1},\dots,p_{e}\in\mathbb{Q}(Z)[T]$ to be irreducible elements such that $\left.P_{m}\right|_{\mu=y_{0},U=Z}=p_{1}\cdots p_{e}$.
@@
By Hilbert’s irreduciblity theorem (see e.g. [10, Chapter 9, Corollary 2.4], which shows that $\mathbb{Q}$ is a Hilbertian field), there are infinitely many $Z_{0}\in\mathbb{Q}$ such that all of the specializations
@@
are irreducible in $\mathbb{Q}[T]$. We note that the set $\bigcup_{i=1}^{e}\{Z_{0}\in\mathbb{Q}\mid\deg(\left.p_{i}\right|_{Z=Z_{0}})\leq 1\}$ is finite from the above observation of the degree of $p_{1},\dots,p_{m}$ in $T$. In addition, for the fixed rational number $y_{0}\neq 1$, the set $\{r\in\mathbb{Q}^{*}\mid r^{n}=y_{0}\text{ for some }n\in\mathbb{Z}\}$ is also finite. As a consequence, there are infinitely many $Z_{0}\in\mathbb{Q}^{*}\setminus\{r\in\mathbb{Q}^{*}\mid r^{n}=y_{0}\text{ for some }n\in\mathbb{Z}\}$ such that $\left.p_{i}\right|_{Z=Z_{0}}\in\mathbb{Q}[T]$ is irreducible and $\deg(\left.p_{i}\right|_{Z=Z_{0}})\geq 2$ for every $i\in\{1,\dots,e\}$, which means that $\left.P_{m}\right|_{\mu=y_{0},U=Z_{0}}$ has no root in $\mathbb{Q}$.
It remains to show the case $m\geq-2$. Let $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ and suppose that $\varphi(L_{K_{0}})$ intersects $\mathbb{R}^{3}$ cleanly along the unknot $K_{1}$. Take a Hamiltonian $H\colon T^{*}\mathbb{R}^{3}\times[0,1]\to\mathbb{R}$ such that $\varphi=\varphi^{1}_{H}$. Consider an involution
@@
$\varphi^{1}_{\overline{H}}(L_{\overline{K_{0}}})$ intersects $\mathbb{R}^{3}=\iota(\mathbb{R}^{3})$ cleanly along $\overline{K_{1}}$. Since $\overline{K_{1}}$ is the unknot and $\overline{K_{0}}$ is the connected sum of the mirror of $K$ and $T_{(2,-2m-1)}$, this contradicts to the result from the previous case. ∎
@@
Then, for any field $\mathbb{k}$,
$\pi_{y,Z}(V_{\mathbb{k}}(K\#4_{1}))\subset\{(y,Z)\in(\mathbb{k}^{*})^{2}\mid\left.P\right|_{\mu=y,U=Z}\text{ has a root in }\mathbb{k}\}.$
@@
If we define $P\in\mathbb{Z}[\mu^{\pm},U][T]$ by (25), then $(U-\mu)\cdot F+\mu Y\cdot G=P(XY)$, which implies that $P(XY)$ is contained in $\mathcal{I}_{B^{\#}}$. Therefore, we have a well-defined unital $R$-algebra map
@@
which maps $T$ to $XY=a_{n,n+3}a_{n+3,n}$.
The rest of argument is parallel to Proposition 3.9. Given a $\mathbb{k}$-valued augmentation $\varepsilon$ of $\mathcal{A}^{\mathrm{knot}}_{*}(K\#4_{1})$, it induces a unital ring homomorphism from $HC^{\mathrm{ab}}_{0}(K\#4_{1})$ to $\mathbb{k}$. By precomposing $J$, we obtain a unital ring homomorphism $R[T]/(P)\to\mathbb{k}$ which sends $\mu$ to $\varepsilon(\mu)$ and $U$ to $\varepsilon(U)$. This means that $\left.P\right|_{\mu=\varepsilon(\mu),U=\varepsilon(U)}\in\mathbb{k}[T]$ has a root in $\mathbb{k}$. Now, the proposition follows from the definition of $V_{\mathbb{k}}(K\#4_{1})$. ∎
@@
Let us abbreviate $\bar{P}=\left.P\right|_{\mu=-1,U=-2}\in\mathbb{Q}[T]$. Note that $(-1,-2)\in\{(y,Z)\mid y\neq Z^{n}\text{ for all }n\in\mathbb{Z}\}$. If we can check that $\bar{P}$ has no root in $\mathbb{Q}$, the theorem immediately follows from Proposition 3.6 and Proposition 3.15.
From the definition of $P$, $\bar{P}=2-\frac{3}{2}T-2T^{3}$. Multiplying by $2$, we obtain a polynomial with integer coefficients $4-3T-4T^{3}$. By the rational root theorem, the only possible rational roots of this polynomial are of the form $\frac{p}{q}$, where $p\in\mathbb{Z}$ is a divisor of $4$ and $q\in\mathbb{Z}$ is a divisor of 4-4. The set of such rational numbers is $\{\pm 1,\pm 2,\pm 4\}\cup\{\pm\frac{1}{2},\pm\frac{1}{4}\}$. We note that $2\cdot\bar{P}(T)=4-T(4T^{2}+3)$. For any $n\in\{\pm 1,\pm 2,\pm 4\}$, the integer $4n^{2}+3$ is not a divisor of $4$, which implies that $2\cdot\bar{P}(n)\neq 0$. One can also check that
@@
Therefore, we can conclude that $\bar{P}$ does not have a root in $\mathbb{Q}$. ∎
@@
For any $\psi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$, we will introduce a $\mathbb{Z}/2$-graded Floer cochain complex
@@
with coefficients in the Laurent polynomial ring $\mathbb{F}_{2}[\lambda^{\pm}]$. It is a slight enhancement of the Floer cochain complex over $\mathbb{F}_{2}$ constructed as in [18, Section 8].
We choose a Hamiltonian $H\colon T^{*}\mathbb{R}^{3}\times[0,1]\to\mathbb{R}$ with compact support such that the Lagrangian submanifold $P_{H}\coloneqq(\varphi^{1}_{H}\circ\psi)(\mathbb{R}^{3})$ intersects $L_{K_{0}}$ transversely. $P_{H}$ is oriented so that the diffeomorphism $\varphi^{1}_{H}\circ\psi\colon\mathbb{R}^{3}\to P_{H}$ preserves orientations. From the orientations of $P_{H}$ and $L_{K_{0}}$, a $\mathbb{Z}/2$-grading $|x|\in\mathbb{Z}/2$ is defined for every $x\in P_{H}\cap L_{K_{0}}$. See [17, Example 2.7, Section 2.d]. Note that the grading is shifted by $1$ if we use the opposite orientation of $P_{H}$. We also choose an almost complex structure $J$ such that $d\lambda_{\mathbb{R}^{3}}(\cdot,J\cdot)$ is a Riemannian metric on $T^{*}\mathbb{R}^{3}$ and $J(\partial_{p_{i}})=\partial_{q_{i}}$ for $i\in\{1,2,3\}$ outside a compact subset of $T^{*}\mathbb{R}^{3}$.